# StackGen: Generating Stable Structures from Silhouettes via Diffusion

### v1

@@ I Introduction
Figure 1: StackGen consists of a diffusion model that takes as inputs a target structure silhouette and a list of available block shapes. The model then generates a set of block poses $\hat{\bm{p}}_{1},\ldots\hat{\bm{p}}_{k}$ that construct a stable structure consistent with the target silhouette. The resulting structure can then be constructed using a robot arm.
@@ III Method
Figure 2: A visualization of StackGen’s transformer-based architecture.
@@ III-A Diffusion Models for SE(3) Block Pose Generation
We adopt denoised diffusion probabilistic models (DDPM) [28] as the core framework of our model. Following DDPM, we start with a forward diffusion process that adds noise to the state space of interest. In our case, given a set of block poses $\bm{p}_{1},\bm{p}_{2},\ldots,\bm{p}_{k}\in\mathbb{R}^{d}$ that compose a stable stack and a diffusion timestep $t\in[1,T]$ that specifies a noise scale. [^1] We normalize poses before applying the diffusion framework. During inference, the generated poses are unnormalized accordingly. Diffusion training with noise-injection follows as:
$\tilde{\bm{p}}^{t}_{i}=\sqrt{\bar{\alpha}_{t}}\bm{p}_{i}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{i},~{}~{}\bm{\epsilon}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $\bar{\alpha}_{t}$ is a coefficient determined by a noise schedule and $\bm{\epsilon}_{i}$ is fresh noise injected in each step. StackGen then provides the noisy poses $\tilde{\bm{p}}^{t}_{i}~{}(i=1,2,\ldots,k)$ to our denoising network $D_{\theta}$ along with the diffusion timestep $t$, shapes $s_{1},\ldots,s_{k}$ and the silhouette of the blocks $S$. This results in the expression
@@
where the notation $X_{1:k}$ is equivalent to $X_{1},\ldots,X_{k}$, and $\bm{\epsilon}_{i}$ is a predicted pose noise for the $i$-th block. With the predicted noises, the training objective for a single sample is
@@
We sample diffusion timestep $t$ uniformly random from $[1,T]$ at each training step.
@@
where $\bm{\epsilon}_{i}$ is given by Eq. 2 and $\bm{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, otherwise $\bm{z}=\bm{0}$. The resulting $\tilde{\bm{p}}^{0}_{1:k}$ are the generated poses.
@@ III-B Model Architecture
Our model (Fig. 2) is built upon the Transformer architecture [35], which can process input tokens that may originate from different modalities. To initialize the process, we use a convolutional neural network (CNN) to predict the block list of a structure from that structure’s silhouette, shown in Figure 1. For training we uniquely encode the number of cubes, rectangles, long rectangles and triangles in a structure using an integer index and proceed by training the CNN $C_{\theta}$ with parametrization $\theta$ to model the joint distribution of block counts, represented by the class probability of each index, using the cross-entropy loss
@@
Given a scene that contains stable stack of $k~{}(\leq N)$ blocks, we extract a list of their poses $\bm{p}_{1},\bm{p}_{2},\ldots,\bm{p}_{k}\in\mathbb{R}^{6}$ and shape embeddings $\bm{s}_{1},\bm{s}_{2},\ldots,\bm{s}_{k}\in\mathbb{R}^{d}$. We use a 6-dimensional pose representation consisting of Cartesian coordinates for translation and exponential coordinates for orientation. The shape embedding is retrieved from a codebook storing unique trainable embeddings for each shape, and the poses are projected into $\mathbb{R}^{d}$ space with an MLP applied independently to each pose. A diffusion timestep $t\in[1,T]$ is also converted to an embedding $\bm{t}\in\mathbb{R}^{d}$. The pose, shape, and diffusion timestep embeddings are summed for each object to obtain $k$ object tokens.
To handle variability in the number of blocks, we make the number of input tokens to the Transformer encoder constant by padding the remaining $N-k$ object tokens with zero vectors, resulting in $N$ tokens independent of $k$.
The silhouette of the block structure is given as a binary image of size $64\times 64$. Following Dosovitskiy et al. [36], we split this into $16$ patches of $16\times 16$ each and use a two-layer MLP to encode each patch independently to obtain silhouette tokens. Sinusoidal positional embeddings [35] are added to the silhouette tokens to retain spatial information.
The object and silhouette tokens are then combined and fed into the Transformer encoder. At the last layer of the encoder, each contextualized block token is projected back to pose space ($\mathbb{R}^{6}$) and supervised with the original noise added to the corresponding pose, following the framework of DDPM. Figure 2 summarizes this process and architecture.
@@ III-C Generating Data
[…]
@@ IV-A Evaluation in simulation
We evaluate our model using a held-out test dataset. Figure 5 shows generated stacks paired with their corresponding silhouettes. In Figure 7, we present a diverse set of stacks produced by the model for a single silhouette, demonstrating its capability for multimodal distribution learning.
Figure 5: Silhouettes from the heldout dataset and rendering of block poses generated by our model.
We aim to evaluate our approach with two metrics: 1) the frequency with which our method generates block poses that form a stable structure; and 2) the consistency of the generated stacks with the target silhouette. The problem of generating a stable structure from a given block list and silhouette can have multiple solutions, so our evaluation technique samples three sets of block poses for each pair of silhouette and block list in the test set. We compare our method against two baselines, the Brute-Force Baseline and the Greedy-Random Baseline.
@@ IV-A1 Brute-Force Baseline
Given a silhouette and a set of available blocks, this algorithm searches for potential placement poses of each block by maximizing a silhouette alignment score given by silhouette intersection while minimizing a collision penalty between predicted blocks. To achieve a high alignment score, for each block we sample 20 coordinates ${(x_{i},0,z_{i})}$ where $x_{i}$ and $z_{i}$ are sampled uniformly from $[-3,3]$ and $\{1,3,5,7\}$, respectively. We perform 20 linear searches from each point along the x-axis in both directions to find poses with optimal alignment and collision measures.
@@ IV-A2 Greedy-Random Baseline
To quantify the diversity of predicted poses, we analyze the predicted results by obtaining a per-layer block list arranged from left to right. Two constructions are considered distinct if their layered block lists differ. The diversity metric is then defined as the number of distinct poses sample poses generated for a given input divided by the total number of samples taken. Figure 7 contains two scenes with average diversity of 83.33%.
TABLE I: Three Side View IoU (%)
Front Side Above Average
Brute-Force Baseline 63.11 58.04 54.49 58.55
Greedy-Random Baseline 58.45 54.50 54.19 55.72
StackGen (ours) 77.03 76.36 70.47 74.62
For stability evaluation, we spawn blocks according to their generated poses, observe their subsequent behavior (i.e., using a simulator for non-real-world experiments), and check whether any of the blocks fall to a layer below where they began, which would lead the sample to be classified as unstable. To assess silhouette consistency, we extract the silhouette of the generated structure after running forward dynamics, compute the intersection over union (IoU) for the silhouettes from three different views (front, side, and top, shown in Table I), and then calculate the average IoU across these views. Unstable (collapsed) structures receive an IoU of zero.
Figure 7: A (left) reference (ground-truth) stack with its silhouette, and (right) a diverse set of structures generated from the silhouette by our model.
TABLE II: Stability and Consistency with Same Diversity
Stability (%) $\uparrow$ IoU (%) $\uparrow$
Brute-Force Baseline 68.13 (1022 / 1500) 58.55
Greedy-Random Baseline 71.93 (1079 / 1500) 55.72
StackGen (Ours) 86.67 (1300 / 1500) 74.62
We evaluated all three models using our pretrained CNN block list predictor on 500 scenes, with three samples generated per scene. Since StackGen achieved a diversity level of $60.47\%$, we set $\sigma=0.6$ in the Greedy-Random Baseline to match this diversity level. As shown in Table II, StackGen significantly outperforms the baselines in both stability and IoU.
IV-B CNN Ablation: Predicted vs. Ground-Truth Block Lists
To ensure that our CNN model did not skew the overall results, we conducted an ablation study following the main experiments. Using 500 scenes from the test dataset, we generated three samples per scene, applying both the CNN-predicted block list $\{\hat{S}_{1:k}\}$ and the ground-truth block list $\{S_{1:k}\}$. As shown in Table III, the quantitative results showed no more than a 2% difference in stability or IoU. These findings confirm that the CNN does not introduce any bottlenecks in our framework. Therefore, all other comparisons are conducted using our CNN predictor.
TABLE III: Ground-Truth vs. Predicted Block List
Stability (%) IoU (%)
StackGen (w/ GT) 88.13 (1322 / 1500) 76.21
StackGen (w/ CNN) 86.67 (1300 / 1500) 74.62
IV-C Block Stacking in the Real World

### v2

@@ I Introduction
Figure 1: StackGen consists of a diffusion model that takes as inputs a target structure silhouette and a list of available block shapes. The model then generates a set of block poses $\{\hat{\bm{p}}_{1},\ldots\hat{\bm{p}}_{k}\}$ that construct a stable structure consistent with the target silhouette. The resulting structure can then be constructed using a robot arm.
@@ III-A Diffusion Models for SE(3) Block Pose Generation
We adopt the denoising diffusion probabilistic model (DDPM) [28] as the core framework underlying StackGen. A probabilistic diffusion model [35] is a generative model comprised of a forward diffusion process and a reverse diffusion process. The forward process is a first-order Markov chain that introduces noise to samples drawn from a data distribution $\bm{p}^{1}\sim q(\bm{p}^{1})$, while the reverse process is a Markov chain that iteratively denoises a noisy input $\bm{p}^{T}\sim\mathcal{N}(0,I)$. The manner by which the model is trained to add and remove noise allows it to generate samples from the target data distribution $q(\bm{p}^{1})$.
In the case of StackGen, $\bm{p}^{1}=\{\bm{p}_{1}^{1},\bm{p}_{2}^{1},\ldots,\bm{p}_{k}^{1}\}$ is the set of object poses, where $\bm{p}_{i}^{1}\in\mathbb{R}^{6}$ is the SE(3) pose of block $i$, and $q(\bm{p}^{1})$ is the distribution over the poses of the blocks that together form a stable stack. [^1] We normalize poses before applying the diffusion framework. During inference, the generated poses are unnormalized accordingly. Following DDPM, StackGen uses a forward diffusion process that injects noise as
$\tilde{\bm{p}}^{t}_{i}=\sqrt{\bar{\alpha}_{t}}\bm{p}_{i}^{1}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{i},~{}~{}\bm{\epsilon}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $t\sim\operatorname{Unif}\{1,2,\ldots,T\}$ is the diffusion timesetep that defines the noise scale, $\bar{\alpha}_{t}$ is a coefficient determined by the noise schedule, and $\bm{\epsilon}_{i}$ is the noise injected in each step. $\bm{p}_{i}^{1}$ is the stable pose of block $i$, $\tilde{\bm{p}}^{t}_{i}$ is noise-injected $p_{i}^{1}$ at timestep $t$. StackGen then provides the noisy poses $\tilde{\bm{p}}^{t}_{i}~{}(i=1,2,\ldots,k)$ to our denoising network (the Transformer in Figure 2) $D_{\theta}$ along with the diffusion timestep $t$, shapes embeddings $\{s_{1},\ldots,s_{k}\}$ and the silhouette of the blocks $S$. This results in the expression
@@
where the notation $X_{1:k}$ is equivalent to $\{X_{1},\ldots,X_{k}\}$, and $\bm{\epsilon}_{i}$ is a predicted pose noise for the $i$-th block. With the predicted noises, the training objective for a single sample is
@@
We sample diffusion timestep $t$ uniformly random from $\{1,2,\ldots,T\}$ at each training step.
@@
where $\bm{\epsilon}_{i}$ is given by Eq. 2 and $\bm{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ if $t>1$, otherwise $\bm{z}=\bm{0}$. The resulting $\tilde{\bm{p}}^{1}_{1:k}$ are the generated stable poses.
Figure 2: A visualization of StackGen’s Transformer-based architecture. We first embed noisy poses $\tilde{\bm{p}}_{1:k}^{t}$, shapes $\bm{s}_{1:k}$ and timestep $t$ to same dimension $d$, then add them together to construct object tokens $\in\mathbb{R}^{k\times d}$. Then we patchify the silhouette and embed them to $\bm{i}_{1:c}$, sum up with the positional encoding $\textrm{Pos}_{1:c}$ to construct silhouette tokens $\in\mathbb{R}^{c\times d}$. Then we feed object, padding and silhouette tokens together to our denoising Transformer $D_{\theta}$ (Eq. 2) in the diagram to predict noise $\hat{\bm{\epsilon}}_{1:k}^{t}$. By running Eq.4 to get next reverse state $\tilde{\bm{p}}_{1:k}^{t-1}$.
@@ III-B Model Architecture
Our model (Fig. 2) is built upon the Transformer architecture [36], which can process input tokens that may originate from different modalities. To initialize the process, we use a convolutional neural network (CNN) to predict the block list of a structure from that structure’s silhouette, shown in Figure 1. The CNN converts the single channel mask input into ten channels, followed by thirty residual layers, max pooling, and two fully connected layers. For training we uniquely encode the number of cubes, rectangles, long rectangles and triangles in a structure using an integer index and proceed by training the CNN $C_{\theta}$ with parametrization $\theta$ to model the joint distribution of block counts, represented by the class probability of each index, using the cross-entropy loss
@@
Given a scene that contains a stable stack of $k$ blocks, we extract a list of their poses $\bm{p}_{1:k}\in\mathbb{R}^{k\times 6}$ and shape embeddings $\bm{s}_{1:k}\in\mathbb{R}^{k\times d}$, where $d=512$ (Fig. 2). For pose $\bm{p_{i}}$, we use a six-dimensional pose representation that consists of Cartesian coordinates for translation and exponential coordinates for orientation. The shape embedding $\bm{s}_{i}$ is retrieved from a codebook that stores unique trainable embeddings for each shape. The poses are projected onto a $d$-dimensional space using an MLP applied independently for each object. Similarly, we project the the diffusion timestep $t\in[1,T]$ onto a $d$-dimensional embedding space. The pose, shape, and diffusion timestep embeddings are summed for each object to obtain $k$ object tokens. In order to handle a variable number of blocks, we use a fixed number of $N$ object tokens to the Transformer encoder and pad the remaining $N-k$ tokens with zero vectors as necessary.
The silhouette $S$ of the block structure is given as a binary image $I$ of size $64\times 64$. Following Dosovitskiy et al. [37], we split the binary image $I$ into $c=16$ patches, each of size $16\times 16$, and independtly encode each patch using a two-layer MLP to obtain silhouette tokens $i_{1:c}\in\mathbb{R}^{c\times d}$. We add sinusoidal positional embeddings [36] $\text{Pos}_{1:c}\in\mathbb{R}^{c\times d}$ to the silhouette tokens to retain spatial information. We then sum the patch embeddings and positional encoding together to get $c$ silhouette tokens.
StackGen then concatenates the object, padding, and silhouette tokens and feeds them into a six-layer Transformer encoder. The Transformer uses a hidden dimension $d=512$ and expands the representation by the feedforward network with dimension $d_{\text{ff}}=2048$. In the self-attention layers, we use $h=8$ self-attention heads, where each head has a dimension of $\frac{d}{h}=64$.
At the last layer of the encoder, StackGen linearly projects each contextualized block token back to pose space ($\mathbb{R}^{6}$) and compute mean squared error (MSE) loss with original noise added to the corresponding pose to conduct supervised learning, following the DDPM framework. Figure 2 summarizes this process and architecture.
[…]
@@ III-C Generating Data
Concretely, we consider a $4\times 4$ grid that serves as a scaffold for block stack designs. We build the initial dense structure from the bottom up, whereby we attempt to place a randomly chosen block (triangles in the top row only) [^2] Throughout this paper we consider four different shapes: {triangle, cube, rectangle, long rectangle}. in the current row without exceeding a maximum width of four. Once at least three cells in a row are occupied, we move on to the next layer. This results in an initial template of a block stack. We then convert the template to a set of corresponding SE(3) poses for the blocks and add a small amount of noise to their horizontal positions. We then use a simulator to verify that the stack is stable under the influence of gravity, render its front silhouette, and add the set of poses along with the silhouette to the dataset. If the stack falls, we simply reject the design. We note that the resulting dataset contains the blocks with slight rotations about the vertical axis, as shown on the right in Figure 3. This is due to the inaccuracy of the physics engine, where the blocks keep slightly sliding and rotating randomly while we run forward dynamics and wait for the other part of the stack to be stable. Although not intended, we keep these in the dataset considering that this randomness helps increase the diversity of the block poses.
@@
Figure 5: A (left) reference (i.e., ground-truth) stack with its silhouette, and (right) a diverse set of structures generated from the silhouette by our model.
@@ IV-A Evaluation in simulation
We evaluate our model using a held-out test dataset. Figure 5 visualizes a diverse set of stacks produced by StackGen for a single silhouette, demonstrating its capability for multimodal distribution learning.
We aim to evaluate our approach with two metrics: 1) the proportion of block configurations generated by our method that are structurally stable; and 2) the consistency of the generated stacks with the target silhouette. The problem of generating a stable structure from a given block list and silhouette can have multiple solutions, so our evaluation technique samples three sets of block poses for each pair of silhouette and block list in the test set. We compare our method against four baselines: two heuristic baselines, the Brute-Force Baseline and the Greedy-Random Baseline, and two learning-based baselines, the Transformer-Regression Baseline and the Transformer-VAE Baseline.
@@ IV-A1 Brute-Force Baseline
Given a silhouette and a set of available blocks, this algorithm searches for potential placement poses of each block by maximizing a silhouette alignment score given by silhouette intersection while minimizing a collision penalty between predicted blocks. To achieve a high alignment score, for each block we sample 20 coordinates ${(x_{i},0,z_{i})}$ where $x_{i}$ and $z_{i}$ are sampled uniformly from $[-3,3]$ and $\{1,3,5,7\}$, respectively. We perform 20 linear searches from each point along the x-axis in both directions to find poses with optimal alignment and collision measures. This algorithm employs a brute-force search over a discretized space of $k\times 400$ candidate positions, which, although not NP-hard, is computationally intensive.
@@ IV-A2 Greedy-Random Baseline
Figure 6: Our learning-based baselines include (left) Tran-VAE: a Transformer-based VAE model with a two-layer encoder and four-layer decoder, and (right) Tran-Reg: a Transformer-based regression model.
IV-A3 Transformer-Regression Baseline
We include a Transformer-based regression model (Fig. 6 (right)) as a non-generative, learning-based method. This baseline provides a means to determine the benefits of using a generative model to capture the intrinsically multimodal distribution of block poses conditioned on the silhouette. To ensure a fair comparison, we train a Transformer with the same number of layers, number of heads, feedforward dimension, and hidden dimension as the Transformer model in StackGen.
IV-A4 Transformer-VAE Baseline
The Tran-VAE baseline (Fig. 6 (left)) employs a Transformer-based variational autoencoder with two encoder layers and four decoder layers, matching the total number of layers (six) in StackGen’s diffusion model. All three learning-based model including StackGen using the same batch size and training epoch to make fair comparison.
The transformer-based baselines (Tran-Reg and Tran-VAE) have positional encodings added to the shape embeddings, whereas StackGen does not. This is because the diffusion models can differentiate different instances of the same block by their distinct noisy poses.
To quantify the diversity of predicted poses, we analyze the predicted results by obtaining a per-layer block list arranged from left to right. Two constructions are considered distinct if their layered block lists differ. The diversity score is then defined as the number of distinct poses sample poses generated for a given input divided by the total number of samples taken. For example, Figure 5 contains two scenes with average diversity score of $\frac{5}{6}=$ 83.33%.
Figure 7: Bar plots that compare StackGen with the baselines in terms of (left) IoU and (right) stability. All methods use the CNN-predicted block list.
@@ IV-A4 Transformer-VAE Baseline
For stability evaluation, we spawn blocks according to their generated poses, observe their subsequent behavior (i.e., using a simulator for non-real-world experiments), and check whether any of the blocks fall to a layer below where they began, which would lead the sample to be classified as unstable. To assess silhouette consistency, we extract the silhouette of the generated structure after running forward dynamics, compute the intersection over union (IoU) for the silhouettes from three different views (front, side, and top, shown in Figure 7 left), and then calculate the average IoU across these views. Unstable (collapsed) structures receive an IoU of zero.
Method Stability Diversity Front IoU Avg IoU
Brute Force 67.33 38.73 62.72 58.45
Brute Force 58.13 39.73 53.39 50.31
Greedy 68.80 56.53 56.01 53.51
Greedy 67.67 57.53 55.13 52.72
Tran-Reg 49.73 38.80 44.57 42.97
Tran-Reg 36.13 46.73 31.78 30.86
Tran-VAE 71.87 56.27 65.46 63.36
Tran-VAE 56.53 60.33 50.19 48.58
StackGen 84.20 66.87 74.39 71.31
StackGen 85.20 64.53 75.08 71.84
TABLE I: Comparison of different methods on stability, diversity, and IoU. All results are presented as percentages (%). The $\dagger$ label indicates models that used the ground-truth block list, those without used the CNN-predicted block list.
We evaluated all five models using our ground-truth and pretrained CNN block list predictor on 500 scenes, with three samples generated per scene. Since StackGen achieved a diversity level of $60.47\%$, we set $\sigma=0.6$ in the Brute-Force and Greedy-Random baselines to match this diversity level. As shown in Table I and Figure 7, StackGen significantly outperforms the baselines in both stability and IoU.
By virtue of being built upon a diffusion model, which is able to represent multimodal distributions, StackGen naturally produces a diverse range of stable designs that meet the silhouette constraints, whereas the heuristic- and learning-based baselines perform noticeably worse in terms of the stability of the resulting stacks, their consistency with the provided sketch (i.e., IoU), and their, perhaps with the exception of the Tran-VAE baseline, their ability to generate diverse designs that match the input silhouette. Meanwhile, although the brute-force method achieves a relatively high IoU score, it requires evaluating an enormous number of pose combinations, resulting in an inference times that exceed six hours (compared to seconds for StackGen). Needless to say, the computational complexity makes them impractical for real-world robotic applications.
IV-B Predicted vs. Ground-Truth Block Lists
As a means of evaluating the influence of our CNN-based prediction of the block list, we compare against methods that are provided access to the ground-truth block list. Using $500$ scenes from the test dataset, we generated three samples per scene and run each model using the CNN-predicted block list $\{\hat{\bm{s}}_{1:k}\}$ and the ground-truth block list $\{\bm{s}_{1:k}\}$. As shown in Table I (models annotated with the $\dagger$ were provided the ground-truth block list), the performance of StackGen in terms of both stability and IoU differs by no more than $2\%$ between the the use of the CNN-predicted and ground-truth block lists, confirming the utility of the CNN-based predictions for StackGen. However, we note the other learning-based methods appear to be highly sensitive to variations in the block list, even though the list is sufficient to generate the stack.
IV-C Block Stacking in the Real World